In the following ordinary annuity, the interest is compounded with each payment, and the payment is made at the end of the compounding period.

Find the amount of time needed for the sinking fund to reach the given accumulated amount. (Round your answer to two decimal places.)

$235 monthly at 5.6% to accumulate $25,000.

The following ordinary annuity, the interest is compounded with each payment, and the payment is made at the end of the compounding period.

Find the accumulated amount of the annuity. (Round your answer to the nearest cent.)

$1500 annually at 6% for 10 years.

In the following ordinary annuity, the interest is compounded with each payment, and the payment is made at the end of the compounding period.

The Oseola McCarty Scholarship Fund at the University of Southern Mississippi was established by a $150,000 gift from an 87-year-old woman who had dropped out of sixth grade and worked for most of her life as a washerwoman. How much would she have had to save each week in a bank account earning 3.9% compounded weekly to have $150,000 after 75 years? (Round your answer to the nearest cent.)

super prize in a contest is $10 million. This prize will be paid out in equal yearly payments over the next 10 years. If the prize money is guaranteed by AAA bonds yielding 6% and is placed into an escrow account when the contest is announced 1 year before the first payment, how much do the contest sponsors have to deposit in the escrow account? (Round your answer to the nearest cent.

.
Answer #1

1

EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+5.6/(12*100))^12-1)*100
Effective Annual Rate% = 5.746
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
25000= 235*(((1+ 5.746/1200)^(n*12)-1)/(5.746/1200))
n(in years) = 7.18

2

EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+6/(12*100))^12-1)*100
Effective Annual Rate% = 6.1678
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 1500*(((1+ 6.1678/100)^10-1)/(6.1678/100))
FV = 19927.69

Please ask remaining parts seperately, questions are unrelated, I have done one bonus